OpenStudy (anonymous):

A person on top of a 19.6 meter tall building throws a ball straight up and it lands on the ground in 5 seconds. What is the initial velocity?

OpenStudy (anonymous):

Using g=10m/s V=21.08m/s

OpenStudy (irishboy123):

x = u*t + (1/2)* a* t* t in this case x= -19.6 t = 5 a = -9.8 You need to memorise the simple equations describing Newtons Laws of motion (constant acceleration) to be fluent with this stuff. Others are v = u + at v^2 = u^2 + 2ax Nothing else to it apart from picking the right equation to use.